OpenStudy (anonymous):

Solve.If the largest of three consecutive even integers is multiplisd by three, the product is 2 less than twice the sum of the first two integers. Find the integers.

OpenStudy (anonymous):

u cant solve the problem ...

OpenStudy (anonymous):

x-1 , x , x+1 given (x+1)(3) = (x-1+x)-2 3x+3 = 2x-1-2 3x+3=2x-3 x=-6 x-1=-7 x+1=-5

OpenStudy (anonymous):

kushashwa, you have three consecutive integers, not three consecutive even integers...

OpenStudy (anonymous):

\[x+2 , x+4,x+6\] \[3x+18 = 2x+6-2\] \[3x+18=2x+4\] \[3x-2x=4-18\] \[x=-14\] \[x+2=-12\] \[x+4=-10\] \[x+6=-8\]

OpenStudy (anonymous):

Number Relation Problem

OpenStudy (anonymous):

got it now jelly delos santos

OpenStudy (anonymous):

x-2, x , x+2 3(x+2)=2(x+x-2)-2 3x+6=2x+2x-4-2 3x+6=4x-6 x=12 so the consecutive even numbers are 10,12,14

OpenStudy (anonymous):

actuallly tomorrow is my exam i am confused doing maths today sorry for the mistakes \[3x+18=2(x+4+x+2)-2\] \[3x+18=4x+12-2\] \[3x+18-10=4x\] \[8=x\] \[x+6=14\] \[x+4=12\] \[x+2=10\]

OpenStudy (anonymous):

what is a true answer when my question?

OpenStudy (anonymous):

the even integers are 10,12,14